# s3: You Don’t Need That Much Data to Train a Search Agent via RL

### v1

@@ Abstract
Pengcheng Jiang, Xueqiang Xu, Jiacheng Lin, Jinfeng Xiao222Independent of co-author’s role at Amazon.2, Zifeng Wang, Jimeng Sun, and Jiawei Han University of Illinois Urbana Champaign †Amazon {pj20,jimeng,hanj}@illinois.edu jfx@amazon.com
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) enables large language models (LLMs) to access and reason over external knowledge by retrieving relevant documents and conditioning generation on them Lewis et al. (2020). As shown in Figure 2, we categorize the evolution of RAG systems into three phases.
@@
Pre-RL-Zero. To improve retrieval quality, subsequent methods enabled more active participation of the LLM during inference. Active RAG techniques Yao et al. (2022); Jiang et al. (2023); Trivedi et al. (2023a) interleaved query generation, retrieval, and reasoning in a multi-turn loop. These systems introduced iterative retrieval but typically relied on zero-shot prompting and lacked trainable components. Self-RAG Asai et al. (2023) distilled such behaviors from larger models into smaller ones via supervised fine-tuning, teaching smaller models to reason and retrieve effectively without external rewards. While these methods improved flexibility and reduced supervision cost, they still did not optimize retrieval using outcome signals.
RL-Zero. The recent emergence of reinforcement learning for retrieval marks a new phase—what we refer to as the RL-Zero era. DeepSeek-R1-Zero Guo et al. (2025) showed that even rule-based, outcome-driven rewards (e.g., answer correctness) can train strong reasoning agents. Building on this idea, DeepRetrieval Jiang et al. (2025) applied RL to train query generators using search-oriented metrics like recall and NDCG. However, these metrics are disconnected from downstream answer quality. Search-R1 Jin et al. (2025) trained a single model to jointly retrieve and generate via reinforcement learning, using exact match (EM) as the reward. While this approach improves answer accuracy, the tight entanglement between search and generation makes it difficult to isolate genuine retrieval improvements (see Figure 3). Moreover, EM is a brittle reward signal—failing to reward semantically correct answers phrased differently.
This motivates a shift toward a modular framework where search and generation are cleanly separated, and optimization focuses purely on search quality with respect to downstream utility Dai et al. (2025). We propose s3, a simple yet powerful framework that trains a search-only agent using a novel reward signal: Gain Beyond RAG (GBR). GBR measures how much better the generator performs when conditioned on retrieved documents from s3, compared to naive top- retrieval. This setup keeps the generator LLM frozen, sidesteps answer token overfitting, and directly optimizes the retrieval component to serve any black-box LLM.
@@ 2.1 Retrieval-Augmented Generation
Large language models (LLMs) have shown impressive generative capabilities Touvron et al. (2023); OpenAI (2023), but their factuality remains bounded Peng et al. (2023) by their training corpora. Retrieval-Augmented Generation (RAG) Lewis et al. (2020); Gao et al. (2023) augments LLMs by prepending retrieved documents to their input, enabling access to up-to-date or domain-specific information. The effectiveness of this setup, however, depends heavily on the retrieval quality. Early efforts improve retrieval through supervised query rewriting Nogueira and Cho (2019); Lin et al. (2023a), where LLMs are fine-tuned to generate better search queries from manually labeled or distilled training data. These methods require significant annotation effort and often optimize for imitation rather than end-task performance. Recent works have introduced Active RAG methods Yao et al. (2022); Trivedi et al. (2023a); Asai et al. (2023); Lyu et al. (2024), which prompt LLMs to iteratively retrieve and reason in a zero-shot or few-shot manner. While flexible, these methods typically rely on handcrafted prompting patterns and lack direct optimization by interacting with environment.
2.2 RL for Agentic Retrieval and Searcher-Centric Optimization
The emergence of reinforcement learning (RL) for large language models has given rise to agentic retrieval, where models interact with search engines and improve by receiving outcome-based feedback—such as whether the final answer is correct. We refer to this shift as the RL-Zero period, sparked by the insight that even simple rewards like answer correctness can elicit strong reasoning and search behavior Guo et al. (2025). Within this paradigm, retrieval-centric methods like DeepRetrieval Jiang et al. (2025) optimize query generation for search metrics (e.g., recall, NDCG), which often fail to reflect answer utility—i.e., whether the retrieved context helps generate a correct answer. Conversely, end-to-end approaches like Search-R1 Jin et al. (2025) train LLMs to retrieve and generate jointly using exact match rewards, but require full model access and entangle search with answer token alignment.
@@ 2.2 RL for Agentic Retrieval and Searcher-Centric Optimization
We introduce s3, a lightweight, model-agnostic framework that equips a tunable search agent with structured, multi-turn access to external knowledge. As illustrated in Figure 4, the searcher LLM interacts with a search engine iteratively: it generates queries, retrieves documents, selects a subset of useful evidence, and decides whether to continue searching. A frozen generator LLM then consumes the accumulated evidence to produce a final answer. To ensure a fair reward baseline, s3 begins by retrieving top- ( in our experiments) documents from the original question, just like naïve RAG. The searcher is trained using the Gain Beyond RAG (GBR) reward, which measures the improvement in generation accuracy when using its retrieved context versus this baseline. This modular design enables targeted optimization of retrieval quality, decoupled from answer generation.
@@ 3.1 Multi-Turn Search-Select Loop
Given a question , the system consists of (1) a searcher LLM (policy) , (2) a search engine , (3) a frozen generator LLM .
s3 first retrieves top- documents using , yielding . A subset is selected to form the initial context. It then performs a sequence of search rounds , structured as follows:
@@
Initialization (Begin with Search). Initializing with ensures the loop begins with the same context as naïve RAG, making the Gain Beyond RAG reward reflect true search improvements.
@@ 3.2 Training via Gain Beyond RAG (GBR)
This reward ensures the searcher is incentivized to retrieve documents that meaningfully enhance the generator’s output quality, independent of surface-form answer similarity. To improve training efficiency, we precompute the baseline accuracy term and restrict training to examples where it equals 0. This effectively filters out questions already solvable by naïve RAG, allowing s3 to focus on harder queries where improved retrieval is essential for generation success.
@@ 3.3 Search Policy Optimization
We optimize the search policy via reinforcement learning using the Gain Beyond RAG (GBR) reward. Each rollout consists of a complete search trajectory: emitted queries, document selections, and a stop decision. Once the final context is constructed, the generator produces an answer, and the GBR reward is computed. The generator remains frozen; gradients are backpropagated only through the search policy. Our method is agnostic to the specific advantage estimation algorithm. In this work, we use Proximal Policy Optimization (PPO) Schulman et al. (2017) due to its strong empirical stability Jiang et al. (2025); Jin et al. (2025). The PPO objective is:
@@
| Corpus: Wikipedia 2018 Karpukhin et al. (2020) | ||||||||
@@
| Corpus: Wikipedia+PubMed+Textbook Xiong et al. (2024) | ||||||||
@@ 4.1 Experimental Setups
Evaluation Metric. We measure performance using Generation Accuracy, which combines a fast span-matching test Ma et al. (2021); Lin et al. (2021) with a lightweight LLM-based correctness check (Figure 13). Given a model prediction and a set of gold answers , we compute:
@@
Datasets. Following prior study Jin et al. (2025), we construct the training set by combining samples from Natural Questions (NQ) and HotpotQA. Since span_check may incorrectly accept answers for questions with semantic negation (e.g., treating “not true” as matching “true”), we remove all yes/no and true/false questions from the training set to ensure reliable reward signals. To focus training on harder examples, we filter out samples where the generator LLM (Qwen2.5-14B-Instruct) already produces a correct answer using naïve RAG retrieval. This reduces the dataset size from 169,615 to 70,286. As later shown in Figure 5, s3 rapidly converges within 15 training steps. For evaluation, we use the checkpoints at step 20. Given a batch size of 120, this corresponds to approximately 2.4k training examples, highlighting the data efficiency of our method. We evaluate on six general-domain QA benchmarks: NQ Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017), PopQA Mallen et al. (2022), HotpotQA Yang et al. (2018), 2WikiMultihopQA Ho et al. (2020), and Musique Trivedi et al. (2022), as well as MIRAGE Xiong et al. (2024), a suite of five medical-domain QA datasets.
Baselines. We compare s3 against a diverse set of RAG systems: (1) End-to-End Fine-tuning. Fully fine-tuned models that jointly retrieve and generate using outcome-based RL or supervision: Search-R1 (3B/7B), SFT (3B), and R1 (7B) where 3B/7B for SFT and R1 are based on Qwen2.5-3B/7B-Instruct. (2) Static Retrieval+Frozen Generator. Methods that retrieve documents using a fixed or scripted strategy, then pass them to a frozen generator: RAG-BM25, RAG-E5: retrieval via BM25 or E5-base Wang et al. (2022). DeepRetrieval-BM25 (3B): RL-trained searcher optimizing recall, paired with BM25. (3) Active Retrieval+Frozen Generator. A diagnostic setting where we extract the documents retrieved during a model’s reasoning trajectory and feed them to a frozen generator: Search-R1-3B/7B (Ret), IRCoT (7B/14B), and Search-o1 (14B) all fall under this category, differing only in whether retrieval is learned (Search-R1) or prompted (IRCoT Trivedi et al. (2023b), Search-o1 Li et al. (2025)). (4) s3. Our s3 trains only the searcher using GBR and forwards selected documents to a frozen generator, with no fine-tuning. We place more details in Appendix A.1.
Models for Training and Evaluation. Throughout all the training processes, we use Qwen2.5-7B-Instruct Yang et al. (2024) as the base searcher LLM to train, and use Qwen2.5-14B-Instruct111In this paper, we use “GPTQ-Int4” version of “Qwen2.5-14B-Instruct” for its high efficiency. We deploy frozen LLMs using vLLM Kwon et al. (2023) for fast inference. as the frozen generator for both answer generation and judge_check for reward computation. For evaluation, we use Claude-3-Haiku as the LLM for judge_check to ensure high evaluation quality. We test three frozen generators: Qwen2.5-7b/14b-Instruct and Claude-3-Haiku. Both training and evaluation are conducted on five NVIDIA A100 80GB PCIe GPUs. RAGEN Wang et al. (2025) and VERL Sheng et al. (2024) are used as the base architecture for multi-turn RL training. We place more details in Appendix A.2.
@@ 5 Results
Compared to IRCoT-14B, which conducts zero-shot retrieval with 2× the parameter count, s3 gains +4.6 points on average. Relative to Search-R1-7B (Ret), which uses the same backbone, s3 improves by +1.5 points while avoiding any generator tuning. These gains are consistent across both single-hop datasets (e.g., 70.0% on NQ) and multi-hop datasets (e.g., 62.4% on HotpotQA), showing that learned search behavior transfers across reasoning complexity.
Medical Domain QA Performance. Table 2 reports performance on the MIRAGE suite Xiong et al. (2024) under both corpus settings. s3 achieves the highest average accuracy (76.6%) when using the combined Wikipedia+PubMed+Textbook corpus, surpassing all retrieval-augmented baselines.
@@
Training Efficiency Table 4 shows that it takes 20 PPO steps (2.4k examples) to train s3, while Search-R1 requires 2,100 steps (170k examples). Even accounting for the higher per-step cost due to LLM-based reward computation, the total wall-clock time is reduced by 33×. Moreover, s3 avoids retriever pretraining and operates with a smaller 7B policy model, making it a practical method for low-resource RL training.
@@
s3 achieves state-of-the-art performance with orders of magnitude less data and compute, suggesting a more sustainable path for RAG optimization.
@@ 6 Conclusion
We present s3, a framework that trains a search-only agent using the Gain Beyond RAG reward. By decoupling search from generation and optimizing only the retriever, s3 outperforms strong baselines with just 2.4k examples. Our results show that targeted search policy learning yields substantial gains in both efficiency and generalization, offering a scalable path for improving RAG systems.
@@ A.1 Baselines Details
IRCoT (7B and 14B). IRCoT222https://github.com/StonyBrookNLP/ircot Trivedi et al. (2023b) is a prompting-based method that alternates between chain-of-thought reasoning and retrieval. It requires no fine-tuning: the model is instructed via prompt to iteratively reason about a question and issue retrieval queries, integrating newly retrieved evidence into its reasoning process. We apply IRCoT using both Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct.
@@
DeepRetrieval-BM25-3B Jiang et al. (2025). This baseline employs a 3B-parameter language model trained with reinforcement learning on retrieval metrics such as recall and NDCG. It learns to generate search queries that maximize the retrieval of relevant documents using a BM25 search engine. Training is conducted on 70k QA examples in NQ dataset with answer span reward (evidence-seeking task in (Jiang et al., 2025)), focusing exclusively on improving retrieval performance, not generation. We use its publicly released checkpoint333https://huggingface.co/DeepRetrieval/DeepRetrieval-NQ-BM25-3B.
@@
Search-R1-3B and Search-R1-7B Jin et al. (2025). These baselines444https://huggingface.co/PeterJinGo/SearchR1-nq_hotpotqa_train-qwen2.5-3b-em-ppo, https://huggingface.co/PeterJinGo/SearchR1-nq_hotpotqa_train-qwen2.5-7b-em-ppo use 3B and 7B parameter models, respectively, and are trained end-to-end to jointly retrieve and generate answers. Reinforcement learning is applied on 170k training examples, using an exact match (EM) reward to guide both retrieval query formulation and answer generation. The model directly integrates search results into its reasoning steps within a single retrieval round.
@@
Search-o1. Search-o1 Li et al. (2025) is an inference-time retrieval controller designed to enhance long-form reasoning in o1-style models such as QwQ and OpenAI’s o1-preview. It is not trained with reinforcement learning or fine-tuned at all. Instead, Search-o1 leverages frozen LLMs and augments them with retrieval by prompting the model to emit search queries mid-reasoning, enclosed in special tokens (e.g., <|begin_search_query|>… Retrieved documents are then post-processed using a Reason-in-Documents module before being injected back into the reasoning flow.
@@
RAG-BM25 and RAG-E5 Lewis et al. (2020). These are naive retrieval-augmented generation baselines with no model training. RAG-BM25 uses top- documents retrieved from a BM25 index, while RAG-E5 retrieves passages using dense retrieval based on E5 embeddings. In both settings, the retrieved documents are prepended to the input prompt and fed into a frozen generator LLM. We set , following prior study Lin et al. (2023b); Jin et al. (2025).
@@
SFT and R1. On general-domain RAG datasets, we train an SFT model with Qwen2.5-3B-Instruct using the same dataset as Search-R1’s 170k NQ+HotpotQA with TRL von Werra et al. (2020) framework. R1 is the “no search” version of Search-R1 Jin et al. (2025), replicating Deepseek-R1-Zero Guo et al. (2025) with a small LLM. We use its publicly released checkpoint555https://huggingface.co/PeterJinGo/R1-nq_hotpotqa_train-qwen2.5-7b-em-ppo-v0.2.
@@
CoT Wei et al. (2022) and Direct Inference. CoT (Chain-of-Thought) prompting instructs the LLM to generate intermediate reasoning steps before producing an answer, without any external retrieval. Direct Inference simply feeds the raw question into the LLM. Neither baseline involves any form of training or finetuning.
@@
To ensure a fair comparison, we set the maximum number of turns to 4 and limit the context to 3 documents per turn for all multi-turn baselines (IRCoT, Search-R1, and Search-o1) and s3, aligning with prior study Jin et al. (2025).

### v2

@@ Abstract
Pengcheng Jiang, Xueqiang Xu, Jiacheng Lin, Jinfeng Xiao222Prior to co-author’s role at Amazon.2, Zifeng Wang, Jimeng Sun, and Jiawei Han University of Illinois Urbana Champaign †Amazon {pj20,jimeng,hanj}@illinois.edu jfx@amazon.com
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) enables large language models (LLMs) to access and reason over external knowledge by retrieving relevant documents and conditioning generation on them lewis2020retrieval. As shown in Figure 2, we categorize the evolution of RAG systems into three phases.
@@
Pre-RLVR. To improve retrieval quality, subsequent methods enabled more active participation of the LLM during inference. Active RAG techniques yao2022react; jiang2023active; trivedi2023interleaving interleaved query generation, retrieval, and reasoning in a multi-turn loop. These systems introduced iterative retrieval but typically relied on zero-shot prompting and lacked trainable components. Self-RAG asai2023self distilled such behaviors from larger models into smaller ones via supervised fine-tuning, teaching smaller models to reason and retrieve effectively without external rewards. While these methods improved flexibility and reduced supervision cost, they still did not optimize retrieval using outcome signals.
RLVR Period. The recent emergence of reinforcement learning with verifiable rewards (RLVR) su2025crossing marks a new phase. DeepSeek-R1-Zero guo2025deepseek showed that even rule-based, outcome-driven rewards (e.g., answer correctness) can train strong reasoning agents. Building on this idea, DeepRetrieval jiang2025deepretrieval applied RL to train query generators using search-oriented metrics like recall and NDCG. However, these metrics are disconnected from downstream answer quality. Search-R1 jin2025search trained a single model to jointly retrieve and generate via reinforcement learning, using exact match (EM) as the reward. While this approach improves answer accuracy, the tight entanglement between search and generation makes it difficult to isolate genuine retrieval improvements (see Figure 3). Moreover, EM is a brittle reward signal—failing to reward semantically correct answers phrased differently.
This motivates a shift toward a modular framework where search and generation are cleanly separated, and optimization focuses purely on search quality with respect to downstream utility dai2025seper. We propose s3, a simple yet powerful framework that trains a search-only agent using a novel reward signal: Gain Beyond RAG (GBR). GBR measures how much better the generator performs when conditioned on retrieved documents from s3, compared to naive top- retrieval. This setup keeps the generator LLM frozen, sidesteps answer token overfitting, and directly optimizes the retrieval component to serve any black-box LLM.
@@ 2.1 Retrieval-Augmented Generation
Large language models (LLMs) have shown impressive generative capabilities touvron2023llama2; openai2023gpt4, but their factuality remains bounded peng2023check by their training corpora. Retrieval-Augmented Generation (RAG) lewis2020retrieval; gao2023retrieval augments LLMs by prepending retrieved documents to their input, enabling access to up-to-date or domain-specific information. The effectiveness of this setup, however, depends heavily on the retrieval quality. Early efforts improve retrieval through supervised query rewriting nogueira2019passage; lin2023train, where LLMs are fine-tuned to generate better search queries from manually labeled or distilled training data. These methods require significant annotation effort and often optimize for imitation rather than end-task performance. Recent works have introduced Active RAG methods yao2022react; trivedi2023interleaving; asai2023self; lyu2024retrieve, which prompt LLMs to iteratively retrieve and reason in a zero-shot or few-shot manner. While flexible, these methods typically rely on handcrafted prompting patterns and lack direct optimization by interacting with environment.
2.2 RL for Agentic Retrieval and Search-Centric Optimization
The emergence of reinforcement learning (RL) for large language models has given rise to agentic retrieval, where models interact with search engines and improve by receiving outcome-based feedback—such as whether the final answer is correct. We refer to this shift as the RL-Zero period, sparked by the insight that even simple rewards like answer correctness can elicit strong reasoning and search behavior guo2025deepseek. Within this paradigm, retrieval-centric methods like DeepRetrieval jiang2025deepretrieval optimize query generation for search metrics (e.g., recall, NDCG), which often fail to reflect answer utility. Conversely, end-to-end approaches like Search-R1 jin2025search train LLMs to retrieve and generate jointly using exact match rewards, but require full model access and entangle search with answer token alignment.
@@ 2.2 RL for Agentic Retrieval and Search-Centric Optimization
We introduce s3, a lightweight, model-agnostic framework that equips a tunable search agent with structured, multi-turn access to external knowledge. As illustrated in Figure 4, the search agent interacts with a search engine iteratively: it generates queries, retrieves documents, selects a subset of useful evidence, and decides whether to continue searching. A frozen generator LLM then consumes the accumulated evidence to produce a final answer. To ensure a fair reward baseline, s3 begins by retrieving top- ( in our experiments) documents from the original question, just like naïve RAG. The search agent is trained using the Gain Beyond RAG (GBR) reward, which measures the improvement in generation accuracy when using its retrieved context versus this baseline. This modular design enables targeted optimization of retrieval quality, decoupled from answer generation.
@@ 3.1 Multi-Turn Search-Select Loop
Given a question , the system consists of (1) a search agent (LLM policy) , (2) a search engine , (3) a frozen generator LLM . s3 first retrieves top- documents using , yielding . A subset is selected to form the initial context. It then performs a sequence of search rounds , structured as:
@@
Begin with Search. Initializing with ensures the loop begins with the same context as naïve RAG, making the Gain Beyond RAG reward (introduced below) reflect real search improvements.
@@ 3.2 Training via Gain Beyond RAG (GBR)
This reward ensures the search agent is incentivized to retrieve documents that meaningfully enhance the generator’s output quality, independent of surface-form answer similarity. To improve training efficiency, we precompute the baseline accuracy term and restrict training to examples where it equals 0. This effectively filters out questions already solvable by naïve RAG, allowing s3 to focus on harder queries where improved retrieval is essential for generation success.
@@ 3.3 Search Policy Optimization
We optimize the search policy via reinforcement learning using the Gain Beyond RAG (GBR) reward. Each rollout consists of a complete search trajectory: emitted queries, document selections, and a stop decision. Once the final context is constructed, the generator produces an answer, and the GBR reward is computed. The generator remains frozen; gradients are backpropagated only through the search policy. Our method is agnostic to the specific advantage estimation algorithm. In this work, we use Proximal Policy Optimization (PPO) schulman2017proximal due to its strong empirical stability jiang2025deepretrieval; jin2025search. The PPO objective is:
@@
| Corpus: Wikipedia 2018 karpukhin2020dense | ||||||||
@@
| Corpus: Wikipedia+PubMed+Textbook xiong2024benchmarking | ||||||||
@@ 4.1 Experimental Setups
Evaluation Metric. We measure performance using Generation Accuracy, which combines a fast span-matching test ma2021replication; lin2021pyserini with a lightweight LLM-based correctness check (Figure 13). Given a model prediction and a set of gold answers , we compute:
@@
Datasets. Following prior study jin2025search, we construct the training set by combining samples from Natural Questions (NQ) and HotpotQA. Since span_check may incorrectly accept answers for questions with semantic negation (e.g., treating “not true” as matching “true”), we remove all yes/no and true/false questions from the training set to ensure reliable reward signals. We evaluate such questions in test data w/ LLMJudge. To focus training on harder examples, we filter out samples where the generator LLM (Qwen2.5-14B-Instruct) already produces a correct answer using naïve RAG retrieval. This reduces the dataset size from 169,615 to 70,286. As later shown in Figure 5, s3 rapidly converges within 15 training steps. For evaluation, we use the checkpoints at step 20. Given a batch size of 120, this corresponds to approximately 2.4k training examples, highlighting the data efficiency of our method. We evaluate on six general-domain QA benchmarks: NQ kwiatkowski2019natural, TriviaQA joshi2017triviaqa, PopQA mallen2023llm_memorization, HotpotQA yang2018hotpotqa, 2WikiMultihopQA xanh2020_2wikimultihop, and Musique trivedi2021musique, as well as MIRAGE xiong2024benchmarking, a suite of five medical-domain QA datasets.
@@
Baselines. We compare s3 against diverse RAG systems: (1) End-to-End Fine-tuning. Fully fine-tuned models that jointly retrieve and generate using outcome-based RL or supervision: Search-R1 (3B/7B), SFT (3B), and R1 (7B) where 3B/7B for SFT and R1 are based on Qwen2.5-3B/7B-Instruct. (2) Static Retrieval+Frozen Generator. Methods that retrieve documents using a fixed or scripted strategy, then pass them to a frozen generator: RAG-BM25, RAG-E5: retrieval via BM25 or E5-base wang2022text. DeepRetrieval-BM25 (3B): RL-trained searcher optimizing recall, paired with BM25. (3) Active Retrieval+Frozen Generator. A diagnostic setting where we extract the documents retrieved during a model’s reasoning trajectory and feed them to a frozen generator: Search-R1-3B/7B (Ret), IRCoT, and Search-o1 all fall under this category, differing only in whether retrieval is learned (Search-R1) or prompted (IRCoT trivedi-etal-2023-interleaving, Search-o1 li2025search). See more details in Appendix A.1.
Models for Training and Evaluation. Throughout all the training processes, we use Qwen2.5-7B-Instruct yang2024qwen2 as the base searcher LLM to train, and use Qwen2.5-14B-Instruct111In this paper, we use “GPTQ-Int4” version of “Qwen2.5-14B-Instruct” for its high efficiency. We deploy frozen LLMs using vLLM kwon2023efficient for fast inference. as the frozen generator for both answer generation and judge_check for reward computation. For evaluation, we use Claude-3-Haiku as the LLM for judge_check to ensure high evaluation quality. We test three frozen generators: Qwen2.5-7b/14b-Instruct and Claude-3-Haiku. Both training and evaluation are conducted on five NVIDIA A100 80GB PCIe GPUs. RAGEN wang2025ragen and VERL sheng2024hybridflow are used as the base architecture for multi-turn RL training. We place more details in Appendix A.2.
@@ 5 Results
Compared to IRCoT-14B, which conducts zero-shot retrieval with 2× the parameter count, s3 gains +4.6 points on average. Relative to Search-R1-7B (Ret), which uses the same backbone, s3 improves by +1.5 points while avoiding any generator tuning. These gains are consistent across both single-hop (e.g., 70.0% on NQ) and multi-hop datasets (e.g., 62.4% on HotpotQA), showing that learned search behavior transfers across reasoning complexity.
Medical Domain QA Performance. Table 2 reports performance on the MIRAGE suite xiong2024benchmarking under both corpus settings. s3 achieves the highest average accuracy (76.6%) when using the combined Wikipedia+PubMed+Textbook corpus, surpassing all retrieval-augmented baselines.
@@
Training Efficiency Table 4 shows that it takes 20 PPO steps (2.4k examples) to train s3, while Search-R1 requires 2,100 steps (170k examples). Even accounting for the higher per-step cost due to LLM-based reward computation, the total wall-clock time is reduced by 33×. Moreover, s3 avoids retriever pretraining and operates with a smaller 7B policy model, making it a practical method for low-resource RL training. s3 achieves state-of-the-art performance with orders of magnitude less data and compute, suggesting a more sustainable path for RAG optimization.
@@ 6 Conclusion
We present s3, a framework that trains a search agent using the Gain Beyond RAG reward. By decoupling search from generation and optimizing only the retriever, s3 outperforms strong baselines with just 2.4k examples. Our results show that targeted search policy learning yields substantial gains in both efficiency and generalization, offering a scalable path for improving RAG systems.
@@ 7 Limitations
Acknowledgements
Research was supported in part by National Science Foundation IIS-19-56151, NSF IIS 25-37827, the Molecule Maker Lab Institute: An AI Research Institutes program supported by NSF under Award No. 2019897, and the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE) by NSF under BRIES Program No. HR0011-24-3-0325.
Contents of Appendix
A. Implementation Details ........................................................................................................................................................................A
A.1 Baselines Details ........................................................................................................................................................................A.1
A.2 Setup Details ........................................................................................................................................................................A.2
A.3 Datasets & Corpora ........................................................................................................................................................................A.3
A.4 Generation Accuracy Computation ........................................................................................................................................................................A.4
A.5 Document Extraction Logic ........................................................................................................................................................................A.5
B. Alignment Study of Evaluation Metrics ........................................................................................................................................................................B
C. Prompts ........................................................................................................................................................................C
D. Scalability Study ........................................................................................................................................................................D
@@ A.1 Baselines Details
IRCoT (7B and 14B). IRCoT222https://github.com/StonyBrookNLP/ircot trivedi-etal-2023-interleaving is a prompting-based method that alternates between chain-of-thought reasoning and retrieval. It requires no fine-tuning: the model is instructed via prompt to iteratively reason about a question and issue retrieval queries, integrating newly retrieved evidence into its reasoning process. We apply IRCoT using both Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct.
@@
DeepRetrieval-BM25-3B jiang2025deepretrieval. This baseline employs a 3B-parameter language model trained with reinforcement learning on retrieval metrics such as recall and NDCG. It learns to generate search queries that maximize the retrieval of relevant documents using a BM25 search engine. Training is conducted on 70k QA examples in NQ dataset with answer span reward (evidence-seeking task in (jiang2025deepretrieval)), focusing exclusively on improving retrieval performance, not generation. We use its publicly released checkpoint333https://huggingface.co/DeepRetrieval/DeepRetrieval-NQ-BM25-3B.
@@
Search-R1-3B and Search-R1-7B jin2025search. These baselines444https://huggingface.co/PeterJinGo/SearchR1-nq_hotpotqa_train-qwen2.5-3b-em-ppo, https://huggingface.co/PeterJinGo/SearchR1-nq_hotpotqa_train-qwen2.5-7b-em-ppo use 3B and 7B parameter models, respectively, and are trained end-to-end to jointly retrieve and generate answers. Reinforcement learning is applied on 170k training examples, using an exact match (EM) reward to guide both retrieval query formulation and answer generation. The model directly integrates search results into its reasoning steps within a single retrieval round.
@@
Search-o1. Search-o1 li2025search is an inference-time retrieval controller designed to enhance long-form reasoning in o1-style models such as QwQ and OpenAI’s o1-preview. It is not trained with reinforcement learning or fine-tuned at all. Instead, Search-o1 leverages frozen LLMs and augments them with retrieval by prompting the model to emit search queries mid-reasoning, enclosed in special tokens (e.g., <|begin_search_query|>… Retrieved documents are then post-processed using a Reason-in-Documents module before being injected back into the reasoning flow.
@@
RAG-BM25 and RAG-E5 lewis2020retrieval. These are naive retrieval-augmented generation baselines with no model training. RAG-BM25 uses top- documents retrieved from a BM25 index, while RAG-E5 retrieves passages using dense retrieval based on E5 embeddings. In both settings, the retrieved documents are prepended to the input prompt and fed into a frozen generator LLM. We set , following prior study lin2023ra; jin2025search.
@@
SFT and R1. On general-domain RAG datasets, we train an SFT model with Qwen2.5-3B-Instruct using the same dataset as Search-R1’s 170k NQ+HotpotQA with TRL vonwerra2022trl framework. R1 is the “no search” version of Search-R1 jin2025search, replicating Deepseek-R1-Zero guo2025deepseek with a small LLM. We use its publicly released checkpoint555https://huggingface.co/PeterJinGo/R1-nq_hotpotqa_train-qwen2.5-7b-em-ppo-v0.2.
@@
CoT wei2022chain and Direct Inference. CoT (Chain-of-Thought) prompting instructs the LLM to generate intermediate reasoning steps before producing an answer, without any external retrieval. Direct Inference simply feeds the raw question into the LLM. Neither baseline involves any form of training or finetuning.
@@
To ensure a fair comparison, we set the maximum number of turns to 4 and limit the context to 3 documents per turn for all multi-turn baselines (IRCoT, Search-R1, and Search-o1) and s3, aligning with prior study jin2025search.